Consider a wave passing through a single slit. What happens to the width of the central maximum of its diffraction pattern as the slit is made half as wide as the first?

A. It becomes one-fourth as wide.

B. It becomes one half as wide

C. Its width does not change

D. It becomes twice as wide.


1
Expert's answer
2021-02-25T11:29:25-0500

The central maximum lies between the first-order minima defined by the relation 


"sin\theta_{dark}=\dfrac{m\lambda}{a}=\dfrac{\lambda}{a}"

Because the angle is small, 

So,

"sin\theta_{dark}\approx tan\theta_{dark}=\dfrac{y_{dark}}{L}"

 

So, the width of the central maximum is proportional to "\dfrac{L\lambda}{a}" .


Thus, the central maximum becomes twice as wide if the slit width a becomes half as wide.

So, according to the question, D option is correct.